OpenStudy (anonymous):

2/5% of what is 0.9?

OpenStudy (anonymous):

5%? or 2/5%

OpenStudy (anonymous):

2/5%

OpenStudy (anonymous):

2/5=0.4 you need just to solve \[(0.4\div100)x=0.9 \implies x=0.9\div(0.4\div100) \implies x=225\]

OpenStudy (anonymous):

Correct !!! I get it . Thanks

OpenStudy (anonymous):

you're welocme

OpenStudy (anonymous):

R U willing to help me out with more problems?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yes sure :)

OpenStudy (anonymous):

finding the percent of change. Round to the nearest tenth of a percent if possible. 1.) The price of a calculator went from $20 last year to $18 this year. What is the percent change this year?

OpenStudy (anonymous):

okay.. you just need to do the following; (18-20)/20=-1/10 then multiply the number by 100% (-1/10)*100%= -10%

OpenStudy (anonymous):

you better write each question in a different page!! :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The price ona pack of colored laser paper went from $3 to $3.60. What is the percent?

OpenStudy (anonymous):

you should try it.. follow the same steps of the previous one: (old price-new price)/old price.. and then multiply by 100%

OpenStudy (anonymous):

ok